OpenStudy (anonymous):

A pyramid has a square base with sides of length s. The height of the pyramid is equal to 1/2 of the length of a side on the base. Which formula represents the volume of the pyramid?

OpenStudy (jdoe0001):

\(\bf \textit{volume of a pyramid}=\cfrac{1}{3}\cdot base\cdot height\) 

OpenStudy (anonymous):

thank you...!!

OpenStudy (jdoe0001):

\(\bf \textit{volume of a pyramid}=\cfrac{1}{3}\cdot {\color{brown}{ base}}\cdot {\color{blue}{ height}} \\ \quad \\ \quad \\ base\ area=s\cdot s\to {\color{brown}{ s^2}}\qquad height={\color{blue}{ \cfrac{1}{2}s}}\)

OpenStudy (jdoe0001):

yw

OpenStudy (anonymous):

thank you..!! am and where are you from?

OpenStudy (jdoe0001):

earth =)

OpenStudy (anonymous):

:).. hehe am gracias eres buen amigo

OpenStudy (jdoe0001):

:)

OpenStudy (anonymous):

speak Spanish..?